### v2

# MindTuner: Cross-Subject Visual Decoding with Visual Fingerprint and Semantic Correction

Abstract
Decoding natural visual scenes from brain activity has flourished, with extensive research in single-subject tasks and, however, less in cross-subject tasks. Reconstructing high-quality images in cross-subject tasks is a challenging problem due to profound individual differences between subjects and the scarcity of data annotation. In this work, we proposed MindTuner for cross-subject visual decoding, which achieves high-quality and rich semantic reconstructions using only 1 hour of fMRI training data benefiting from the phenomena of visual fingerprint in the human visual system and a novel fMRI-to-text alignment paradigm. Firstly, we pre-train a multi-subject model among 7 subjects and fine-tune it with scarce data on new subjects, where LoRAs with Skip-LoRAs are utilized to learn the visual fingerprint. Then, we take the image modality as the intermediate pivot modality to achieve fMRI-to-text alignment, which achieves impressive fMRI-to-text retrieval performance and corrects fMRI-to-image reconstruction with fine-tuned semantics. The results of both qualitative and quantitative analyses demonstrate that MindTuner surpasses state-of-the-art cross-subject visual decoding models on the Natural Scenes Dataset (NSD), whether using training data of 1 hour or 40 hours.
Introduction
Do our brains form unified perceptions when we observe similar objects? Do our unique understandings influence these perceptions differently? The human brain exhibits substantial anatomical similarities in terms of functional organization, including shared attributes like memory, functional connectivity, and visual cortex functions (Chen et al. 2017; Fingelkurts, Fingelkurts, and Kähkönen 2005; Stringer et al. 2019). However, individual neural biases always exist due to inherent differences (Wang, Murai, and Whitney 2020). Understanding both the similarities and gaps in perception has profound implications for the fields of Artificial Intelligence (AI) (Nie et al. 2023; Zhao et al. 2014) and Brain-Computer Interface (BCI) research (Wang et al. 2009). Visual decoding is a straightforward way to understand the brain, where functional magnetic resonance imaging (fMRI) is a widely embraced non-invasive tool used to decode natural visual stimuli, revealing intricate perceptual and semantic details in the cerebral cortex (Linden 2021). Consequently, fMRI has garnered considerable attention in image retrieval and reconstruction tasks.
Open-source large-scale fMRI datasets, such as the Natural Scenes Dataset (NSD) (Allen et al. 2022), advance deep learning models to shine in fMRI decoding. Pre-trained cross-modality models like CLIP (Radford et al. 2021) and Stable Diffusion (Rombach et al. 2022) offer effective representation space and models for high-quality visual reconstruction. A large body of literature demonstrates the feasibility of training single-subject decoding models to reconstruct high-fidelity images (Lin, Sprague, and Singh 2022; Scotti et al. 2023; Chen et al. 2023; Takagi and Nishimoto 2023; Lu et al. 2023). However, single-subject decoding has drawbacks, including the need to train a unique model for each subject, making it challenging to generalize to new subjects and requiring a substantial amount of fMRI data for training. As is widely recognized, acquiring a large amount of fMRI data for each subject is time-consuming, labor-intensive, and impractical in practical scenarios. Unfortunately, most current research focuses on single-subject visual decoding rather than exploring the challenging commonalities of the human brain. Consequently, there is an urgent need for cross-subject decoding models that can be effectively transferred to new subjects and perform well in the few-shot setting, as depicted in Figure LABEL:1_hour.
The key to cross-subject few-shot decoding lies in effectively utilizing extensive prior knowledge from other subjects or additional modalities. On the one hand, one successful strategy for leveraging knowledge from other subjects involves aligning them to a shared space. Ridge regression is commonly employed for this purpose, aligning voxel inputs from different subjects (Scotti et al. 2024; Ferrante, Boccato, and Toschi 2023). This approach is preferred due to the low signal-to-noise ratio in fMRI data, where complex non-linear models tend to overfit noise. Nonetheless, the process of visual information perception and generating brain activity in each individual incorporates unique components, referred to as the visual fingerprint (Wang, Murai, and Whitney 2020) (refer to Preliminary for more detailed analysis). Current linear alignment methods only enable new subjects to conform to the shared components across subjects, neglecting the perception difference derived from their distinctive visual fingerprint and resulting in limited performance.
On the other hand, an additional strategy involves leveraging multimodal data. Previous alignment methods focused solely on the visual modality, as a means of adapting to the inputs of Stable Diffusion. However, this alignment approach is susceptible to slight disturbances, leading to semantic errors in the generated images. For visual decoding tasks, the textual modality is highly relevant and is verified effective in enhancing visual decoding semantically (Scotti et al. 2024). However, previous approaches incorporating text have placed excessive emphasis on directly aligning fMRI with detailed textual descriptions to facilitate the reconstruction process (Takagi and Nishimoto 2023), which intuitively lacks rationality and yields poor performance. Considering that subjects in visual stimulation experiments lack direct interaction with the textual modality and that individual understanding of visual stimuli varies, the relationship between fMRI and text should be considered implicit.
In this paper, we propose MindTuner: a cross-subject visual decoding framework. Inspired by visual fingerprint in Brain Science (Wang, Murai, and Whitney 2020), and with the help of Low-Rank Adaptation (LoRA) for large model lightweight fine-tuning. In correspondence to the above two strategies, we first propose the combination of non-linear Skip-LoRAs and LoRAs to learn the visual fingerprint of new subjects, which are injected into the fMRI encoding network to correct visual perception difference. In addition, we design a Pivot module that uses images as the central modality to bridge fMRI and text. The Pivot helps to correct the reconstructed image with fine-tuned semantics. Our contributions are summarized as follows:
  • MindTuner makes the first attempt to introduce LoRA as a subject-level fine-tuning module in cross-subject decoding and further elaborately design non-linear Skip-LoRA. Their combination shows excellent capability to learn subjects’ visual fingerprint.
  • We introduce a novel fMRI-to-text retrieval paradigm with a Pivot using the image modality. The Pivot conducts semantic correction with label prompts to enhance fMRI-to-image reconstruction.
  • We evaluate our method on the NSD dataset, and it establishes SOTA decoding performance whether using training data of 40 hours (full) or 1 hour(2.5% of full).
Related Work
Cross-Subject Functional Alignment
In visual decoding, single-subject models have exposed the issue of excessive reliance on the data volume of individual subjects, leading researchers to shift toward cross-subject studies. Functional alignment of different brains is considered to be more effective than anatomical alignment. Previous functional alignment methods were mainly divided into two perspectives: fMRI data itself and downstream tasks. Among a series of methods starting from the fMRI perspective, Bazeille et al. (Bazeille et al. 2019) and Thual et al. (Thual et al. 2023) minimize an optimal transport cost between voxels of different brains. The methods based on fMRI self-supervision (Chen et al. 2023; Qian et al. 2023) emphasize obtaining common latent representations between subjects through autoencoders. Wills Aligner  (Bao et al. 2024) aligns data from different subjects using the fsaverage template. On another cross-subject alignment research path, most of these methods focus on the shared knowledge among subjects and overlook the complex nonlinear relationships between subjects due to concerns about overfitting. Linear fitting  (Ferrante, Boccato, and Toschi 2023) was performed on the responses of subjects to common images, achieving high-quality cross-subject visual reconstruction, but requiring subjects to see the same images. Mindeye2  (Scotti et al. 2024), MindBridge  (Wang et al. 2024) and UMBRAE  (Xia et al. 2024b) improved this and achieved impressive results, but still losing subject-specific features and ignoring the more complex relationships between subjects, which is not flexible.
Text Modality in Visual Decoding
Decoding visual stimuli from fMRI has been a long-standing endeavor, primarily focusing on the image modality  (Takagi and Nishimoto 2023; Lu et al. 2023; Gong et al. 2024b; Ozcelik and VanRullen 2023; Scotti et al. 2023). However, an increasing number of studies have highlighted the role of text. UniBrain (Mai and Zhang 2023) directly aligns fMRI to the text representation via ridge regression and then completes the brain caption task with the text generation model. MindEye2 (Scotti et al. 2024) obtains a better caption by placing the generated image representation into the image captioning model, which is used to smooth the generated image. NeuroClips utilizes the pretrained BLIP-2  (Li et al. 2023) model to achieve text-assisted video reconstruction  (Gong et al. 2024a). However, these methods either do direct fMRI-to-text alignment or direct image-text alignment, ignoring the indirect relationship between fMRI and text from the perspective of intuitive understanding.
Preliminary on visual fingerprint
Visual stimuli are processed by the Human Visual System(HVS). Therefore, as stated in previous research (Xia et al. 2024a), brain responses such as fMRI are closely linked to our visual system. Although presented in a systematic manner, there are still significant differences in the visual systems of different individuals. Research of visual fingerprint emphasizes that idiosyncratic biases exist in the underlying representation of visual space propagate across varying levels of visual processing (Wang, Murai, and Whitney 2020). Using a position-matching task will find stable subject-specific compressions and expansions within local regions throughout the visual field. As shown on the left of Figure 1, in experiments, subjects were fixated at the center, and a target was displayed briefly at one of five possible eccentricities (depicted by dashed lines, which were not visible in the experiment). After the target disappeared, subjects moved the cursor to match the target’s location. Linear models were used to fit the Distortion Indices(DI), which measured the degree of spatial distortion between subjects:
$DI_{self}$ $\sim\beta_{0}+\beta_{1}\times self,$ (1)
$DI_{others}$ $\sim\beta^{\prime}_{0}+\beta^{\prime}_{1}\times others.$
As shown on the right of Figure 1, the experiment results show that subjects have their own visual fingerprint, with both linear and non-linear components. Within-subject similarity (r=0.71) is significantly higher than between-subject similarity (r=0.22), suggesting that each individual subject has their own unique spatial distortions that are consistent within themselves and distinguished from others.
Figure 1: Visual fingerprint experiments across subjects. ’Within’ denotes Pearson correlation coefficient of Distortion Indices in within-subject experiments, while ’between’ denotes between-subject.
Method
Figure 2: Schematic diagram of MindTuner. The training process was split into two phases: Multi-subject Pre-training and Cross-subject Fine-tuning, in which the corresponding modules were trained. The predicted embeddings are first obtained through MindTuner, and then the preliminary reconstructed image was obtained by SDXL unCLIP. The final reconstructed image is obtained by text retrieval and semantic correction by SDXL Image-Variation.
The task of cross-subject visual decoding consists of two parts: multi-subject pre-training and new-subject fine-tuning. Given a neural dataset with brain activities of subject $s$, it involves the reconstruction of visual images $I\in\mathbb{R}^{3\times H\times W}$ through a core pipeline whose inputs are flattened and aligned fMRI voxels $V\in\mathbb{R}^{1\times d_{s}}$ after ROI extraction, where $d_{s}$ denotes voxel numbers. To simplify notation, each $(I,V)$ denotes data from the original subjects for this section. Our goal is to optimize the $\mathcal{F(\cdot)}$, so that $\mathcal{F}(V)=\hat{I}$, where $\hat{I}$ best approximates $I$. New-subject fine-tuning adheres to the same path, albeit with scarce data. For the multi-subject pre-training, we follow the pipeline of MindEye2, while for the new-subject fine-tuning, we plug into visual fingerprint and Pivot as shown in Figure 2.
Multi-Subject Pre-training
Multi-Subject Functional Alignment
It should be noted that the brain structures of different subjects vary, as do the number of voxels obtained. Consequently, a mapping model is required to align the voxel inputs from different subjects to the same dimension as the inputs to the model in the phase of multi-subject pre-training. Here, we employ linear function alignment to learn shared-subject fMRI latent space $M\in\mathbb{R}^{1\times d_{0}}$($d_{0}$ denotes shared input dimension). This is achieved through subject-specific ridge regression, as detailed below:
$M=\textbf{Ridge}^{(s)}(V)$ (2)
MLP Backbone
In order to achieve high-fidelity reconstruction, it is necessary to utilize a substantial quantity of CLIP image embedding. The OpenCLIP ViT-bigG/14 space is employed for alignment, with an image embedding dimension of $256\times 1664$. Mapped inputs $M$ are fed into an MLP backbone comprising four residual blocks and a tokenization layer which transforms the input from a dimension of $d_{0}$ to $256\times 1664$. The MLP Backbone $\epsilon(\cdot)$ serves to convert the fMRI to the intermediate backbone embedding space: $Z=\epsilon(M)$. It should be noted that all subjects shared the same MLP backbone following multi-subject functional alignment. Subsequently, the backbone embeddings are conveyed into three submodules for retrieval, high-level reconstruction, and low-level reconstruction.
Retrieval Submodules
A straightforward approach to performing the retrieval task is to conduct a shallow mapping of the backbone embeddings and supervise it with an fMRI-to-image CLIP contrastive loss. In the case of limited fMRI data, the application of appropriate data augmentation techniques can facilitate the convergence of the model. A recently proposed voxel mixture paradigm, based on MixCo  (Kim et al. 2020), has demonstrated effectiveness. Two raw fMRI voxels $V_{i}$ and $V_{j}$ are mixed into $V_{mix_{i,j}}$ using a factor $\lambda$ sampled from the Beta distribution:
$V_{mix_{i,j}}$ $=\lambda_{i}V_{i}+(1-\lambda_{i})V_{j},$ (3)
$M_{mix_{i,j}}$ $=\textbf{Ridge}^{(s)}(V_{mix_{i,j}}),$
$Z_{mix_{i,j}}$ $=\epsilon(M_{mix_{i,j}})$
where $j$ denotes an arbitrary mixing index in the batch. The forward mixed contrastive loss MixCo is formulated as:
$\mathcal{L}_{MixCo}=-\frac{1}{|B|}\sum_{i=1}^{|B|}[\lambda_{i}\log\frac{\exp(Z_{mix_{i,j}}\cdot\ g_{i}/{\tau})}{\sum_{m=1}^{|B|}\exp(Z_{mix_{i,j}}\cdot\ g_{m}/{\tau})}$ (4)
$+(1-\lambda_{i})\log\frac{exp(Z_{mix_{i,j}}\cdot\ g_{j}/{\tau})}{\sum_{m=1}^{|B|}\exp(Z_{mix_{i,j}}\cdot g_{m}/{\tau})}]$
where $g$ denotes ground-truth CLIP image embeddings, $\tau$ denotes a temperature hyperparameter, and $B$ is the batch size. Here we use the bidirectional loss $\mathcal{L}_{BiMixCo}$.
Low-level and High-level Submodules
The low-level pipeline is a widely utilized technique for the enhancement of low-level visual metrics in reconstruction images. This involves the mapping of voxels to the latent space of Stable Diffusion’s Variational AutoEncoder (SDVAE), which serves as a surrogate for the reconstruction. The pipeline comprises an MLP and a CNN upsampler with L1 loss in Stable Diffusion’s latent embeddings $z$.
$\mathcal{L}_{lowlevel}=\frac{1}{|B|}\sum_{i=1}^{|B|}|z_{i}-\hat{z_{i}}|$ (5)
Conversely, a high-level pipeline places greater emphasis on semantic alignment. Inspired by DALLE·2 (Ramesh et al. 2022), a diffusion prior is recognized as an effective means of transforming backbone embeddings into CLIP ViT image embeddings, in which mean square error loss is used (further insights on the deployment of diffusion priors can be found in Appendix B.4):
$\mathcal{L}_{prior}=\frac{1}{|B|}\sum_{i=1}^{|B|}||g_{i}-\hat{g_{i}}||^{2}_{2}$ (6)
Thus, the end-to-end loss for multi-subject pre-training is:
$\mathcal{L}_{multi}=\mathcal{L}_{prior}+\alpha_{1}\mathcal{L}_{lowlevel}+\alpha_{2}\mathcal{L}_{BiMixCo}$ (7)
Where $\alpha$ denotes the weight to balance multiple losses.
New-Subject Fine-tuning
Low-Rank Adaptation
Previous work has demonstrated the effectiveness of low-rank adaptation in fine-tuning large language models with significantly fewer trainable parameters. It is well suited to our multi-subject decoding task for two reasons. First, most of the current mainstream models for fMRI decoding are MLP-based models that contain a large number of linear layers, whereas LoRA has been shown to achieve good fine-tuning results in the linear layers. Second, in cross-subject scenarios, fMRI data from new subjects are usually scarce, and full-tuning the whole model is usually difficult to grasp, leading to a certain degree of overfitting. For each pre-trained weight matrix $\in\mathbb{R}^{d_{in}\times d_{out}}$ in the multi-subject model, where $d_{in}$ denotes input dimension and $d_{out}$ denotes output dimension, gradient update is constrained with a low-rank decomposition for new-subject adapter matrix $\Delta W$:
$W+\Delta W=W+BA$ (8)
where $W$ is keeped frozen, $B\in\mathbb{R}^{d_{out}\times r}$, $A\in\mathbb{R}^{r\times d_{in}}$, $r$ is the rank and $r\ll min(d_{in},d_{out})$. At the beginning of the training phase, the parameters of the matrix $B$ are randomly initialized, and $A$ is initialized to zero, which ensures that the initial output of the LoRA block is all zeros.
non-linear Skip-LoRAs
[…]
$M_{0}=\textbf{Ridge}^{(s)}(V),$ (9)
$M_{h}=M_{h-skip}+M_{h-linear},$
$M_{h-linear}=\epsilon_{h}(M_{h-1})+\textbf{LoRA}_{h}(M_{h-1}),$
$M_{h-skip}=\textbf{Skip-LoRA}_{h}(V)$
Here, we use the Pearson correlation coefficient to define the non-linear correlation loss:
$\mathcal{L}_{skip}=\frac{1}{|B|}\frac{1}{|H|}\sum_{i=1}^{|B|}\sum_{h=1}^{|H|}|Pearson(M_{h-linear},M_{h-skip})|$ (10)
Pivot with Adaptive Projector
In addition to pixel-level alignment, semantic-level matching is equally important for reconstructing the semantic integrity of an image. In visual decoding tasks, fMRI is directly associated with the image stimulus, and semantic information is implicit in fMRI. More details on the Adaptive Projector can be found in Appendix B.3. For new subjects, we add the additional image-text loss to the model to constrain the semantic information:
$\mathcal{L}_{pivot}=-\frac{1}{|B|}\sum_{i=1}^{|B|}\log\frac{\exp(p_{i}\cdot\ t_{i}/{\tau})}{\sum_{m=1}^{|B|}\exp(p_{i}\cdot\ t_{m}/{\tau})}$ (11)
where $p$ denotes the projector’s output from image tokens $g$, and $t$ denotes the CLS embeddings (CLS denotes the classification token) of paired text. During new-subject fine-tuning, the adaptive projector is trainable. Thus, the end-to-end loss for new-subject fine-tuning is:
$\mathcal{L}_{new}=\mathcal{L}_{multi}+\alpha_{3}\mathcal{L}_{skip}+\alpha_{4}\mathcal{L}_{pivot}$ (12)
Semantic Correction
MindEye2 has found that the reconstructed images from SDXL unCLIP after token alignment have fuzzy semantics, so the aligned embeddings are fed into the image captioning model GIT (Wang et al. 2022) to get the semantics for refinement. We have found that single category words are sufficient to accomplish the refinement task. In addition, the captions generated by MindEye2 are completely dependent on aligned embeddings, which can only make the image semantics clearer and cannot correct the semantics of the reconstructed images, as shown in Figure 3. Taking advantage of our adaptive projector, we can define category description as an fMRI-to-text retrieval task using a simple text prompt $[class]$. In this way, we achieve more accurate category reconstruction.
Figure 3: SDXL unCLIP reconstructions and SDXL Image-Variation by MindEye2’s refinement or our correction.
Method Low-Level High-Level Retrieval
[…]
Takagi… 0.246 0.410 78.9% 85.6% 83.8% 82.1% 0.811 0.504 - -
Ozcelik… 0.273 0.365 94.4% 96.6% 91.3% 90.9% 0.728 0.422 18.8% 26.3%
MindEye1 0.319 0.360 92.8% 96.9% 94.6% 93.3% 0.648 0.377 90.0% 84.1%
MindEye2 0.322 0.431 96.1% 98.6% 95.4% 93.0% 0.619 0.344 98.8% 98.3%
MindTuner(Ours) 0.322 0.421 95.8% 98.8% 95.6% 93.8% 0.612 0.340 98.9% 98.3%
MindEye2(1 hour) 0.195 0.419 84.2% 90.6% 81.2% 79.2% 0.810 0.468 79.0% 57.4%
MindTuner(1 hour) 0.224 0.420 87.8% 93.6% 84.8% 83.5% 0.780 0.440 83.1% 76.0%
Table 1: Quantitative comparison of MindTuner’s performance against other methods. All results are averaged across subjects 1, 2, 5, and 7 from the Natural Scenes Dataset. The results of other methods are taken from MindEye2(Scotti et al. 2024), using either 40-hour data or 1-hour data. Some methods were not included in the comparison because they used past 37-hour data or were not open-source. Missing values occur when metrics are not applicable. Bold font signifies the best performance, while underlined text indicates the second-best performance (See Appendix B.1 for more details about the metrics).
Experiment
Datasets
Natural Scenes Dataset (NSD) [^1] https://naturalscenesdataset.org (Allen et al. 2022) is an extensive 7T fMRI dataset gathered from 8 subjects viewing images from the MSCOCO-2017 dataset, which contains images of complex natural scenes. Participants viewed three repetitions of 10,000 images with a 7-Tesla fMRI scanner over 30–40 sessions, with one session including 750 fMRI trials lasting for 1 hour. Details of the dataset information can be found in Appendix A. In this paper, we conducted 4 types of experiments, image retrieval and reconstruction in the next section, and text retrieval and brain correlation experiments in Appendix C.
Implementation details
All of our fine-tuning experiments were run for 150 epochs on two Tesla v100 32GB GPUs with a batch size of 10. The experimental parameter settings for 1-hour and 40-hour data are consistent. For fine-tuning experiments, the loss weight is set to $[\alpha_{1},\alpha_{2},\alpha_{3},\alpha_{4}]=[0.5,1.0,1.5,0.5]$ and the rank $r$ is set to 8 for all LoRA blocks, including Skip-LoRAs. We use the AdamW (Loshchilov and Hutter 2017) for optimization, with a learning rate set to 3e-4, to which the OneCircle learning rate schedule (Smith and Topin 2019) was set. During the training process, we use data augmentation from images and blurry images and replace the BiMixCo with SoftCLIP (Scotti et al. 2023) loss in one-third of the training phase. In the inference stage, we only generated a reconstructed image once and did not make multiple selections. The final high-level reconstructed image and low-level blurry image are simply weighted and averaged in a ratio of 3:1. In the inference stage of semantic correction, we use 80 category nouns from MSCOCO as correction texts and classify them in the form of text retrieval.
Results and Analysis
Image and Brain Retrieval
Image retrieval refers to retrieving the image embeddings with the highest cosine similarity based on fMRI embeddings on the test set. In Table 1, there is only a slight improvement in retrieval accuracy within 40 hours of data. This can be attributed to the fact that the retrieval accuracy is already close to the upper limit (only about 1% short of reaching 100%), which is affected by the noise of the fMRI dataset itself. However, when only 1 hour of data was available, MindTuner’s retrieval accuracy was significantly higher than MineEye2, 4.1% higher for image retrieval and 18.6% higher for brain retrieval. As MindEye2 benefits from multi-subject pre-training and new-subject fine-tuning with linear heads, this suggests that the visual fingerprint we introduced significantly improves the performance of the model when fMRI data are scarce.
Figure 4: MindTuner vs MindEye2 reconstructions from fMRI brain activity with only 1 hour of data.
Figure 5: MindTuner vs MindEye2 reconstructions from fMRI brain activity with 40 hours of data.
Method Trainable Low-Level High-Level Retrieval
[…]
MindEye2 64.4M(-) 0.235 0.428 88.0% 93.3% 83.6% 80.8% 0.798 0.459 94.0% 77.6%
with Adaptive Projector 66.5M(+2.1M) 0.233 0.426 87.8% 93.0% 84.0% 81.2% 0.794 0.454 93.8% 77.3%
with only LoRAs 74.6M(+9.4M) 0.261 0.427 90.3% 94.2% 84.8% 84.0% 0.784 0.441 93.7% 85.6%
LoRAs+Skip-LoRAs 74.6M(+10.2M) 0.264 0.427 90.8% 94.8% 85.1% 84.2% 0.780 0.437 94.5% 87.7%
MindTuner(non-linear) 76.7M(+12.3M) 0.191 0.383 82.7% 88.0% 77.3% 76.0% 0.848 0.492 27.4% 56.7%
MindTuner 76.7M(+12.3M) 0.262 0.422 90.6% 94.9% 85.8% 84.6% 0.774 0.433 94.2% 87.4%
Table 2: Ablations on the modules of MindTuner. Note that all model parameters for MindEye2 are 2.2B, and we only reported the trainable parameters required for new-subject fine-tuning. MinTuner(non-linear) denotes adding an activation function to the ridge regression head.
$rank$ Trainable Skip-LoRAs Low-Level High-Level Retrieval
[…]
$4$ 71.6M(+7.2M) 0.4M 0.263 0.422 90.4% 94.9% 84.8% 84.6% 0.778 0.436 94.4% 87.9%
$8$ 76.7M(+12.3M) 0.8M 0.262 0.422 90.6% 94.9% 85.8% 84.6% 0.774 0.433 94.2% 87.4%
$16$ 86.7M(+22.3M) 1.6M 0.262 0.422 90.5% 95.1% 85.2% 85.2% 0.776 0.434 93.7% 87.2%
Table 3: Ablations on the rank $r$ of MindTuner.
Image Reconstruction
Image reconstruction aims to restore the original image as seen by the subjects, and two levels of evaluation metrics assess the quality of the reconstructions. Previous research has improved the performance of these above metrics in single-subject models by various means. The cross-subject task discussed in this paper tests the ability of the model to exploit the commonalities of multi-subject and to migrate new subjects, ultimately realizing the goal of using less fMRI data for few-shot learning. Here, we report the reconstruction results of MindTuner using 1-hour and 40-hour training data. Table 1 reflects the quantized performance comparisons, under two different training data sizes. It can be seen that when using the full NSD dataset, MindTuner achieves better performance on high-level metrics; when only 1 hour of data is available for training, MindTuner outperforms MindEye2 on all metrics. The visualization results of 1-hour and 40-hour reconstructions can be seen in Figure 4 and Figure 5. It can be seen that the quality of the reconstructed images at 40 hours is significantly higher than at 1 hour as the training data increases. Meanwhile, our MindTuner is better than MindEye2 in both semantic completeness and category accuracy, demonstrating the superiority of the overall model. Our visualization excludes the other three methods in Table 1 because the image semantics generated by these three methods are very unclear. Interestingly, we also found that images generated directly from SDXL unCLIP outperformed MindTuner’s corrected images at a high level, albeit visibly distorted. Further results are in Appendix D.1.
Ablations
In this section, we explored the effectiveness of each component of our method through ablation experiments. All results were pre-trained on Subjects 2-8, and fine-tuning was performed on Subject 1 using 1 hour of data.
MindTuner’s modules. Further experiments were conducted to assess the efficacy of each module. As can be observed in Table 2, the incorporation of LoRAs markedly enhances the model’s capacity. The incorporation of Skip-LoRAs resulted in a convergence of all image reconstruction metrics towards the performance of the complete model, with the retrieval accuracy even exceeding that of the complete model. This can be attributed to the balance of additional image-to-text aligned losses, which has the effect of slightly reducing the performance of the retrieval submodule in the complete model. Furthermore, the incorporation of an adaptive projector has enhanced the performance of high-level reconstruction, and additional visualizations in Figure 3 have also demonstrated the efficacy of semantic correction. However, this approach resulted in the compromise of certain low-level effects. In comparison to the 2.2B multi-subject pre-trained main model, MindTuner resulted in a mere 12.3M increase in parameters (0.56%), yet achieved superior outcomes. Furthermore, an activation function was added to the ridge head to make the entire alignment head non-linear. However, this resulted in a serious overfitting phenomenon, with all metrics exhibiting a significant decline, which indicates that the Skip-LoRAs in MindTuner offer a more suitable alternative for non-linear fMRI alignment of diverse subjects.
The rank $r$ of Skip-LoRAs and LoRAs. Further experiments were conducted on the most critical variable in the LoRAs, namely the rank $r$. As can be observed in Table 3, when the rank $r=4/16$, the overall ability of the model does not show significant changes. This finding is consistent with the results presented in the LoRA paper (Hu et al. 2021), which highlighted that when r is within the range of 2 to 16, the fine-tuning effect remains unchanged. Furthermore, no substantial evidence of overfitting was identified. This can be attributed to the fact that the parameters in Skip-LoRAs have remained relatively modest, below 2M.
Neuroscience interpretability
To investigate the interpretability of MindTuner in neuroscience, we conduct experiments to explore where the subject’s non-linear relationship comes from in the visual cortex. We utilize pycortex (Gao et al. 2015) to project the weights of each voxel in the first layer of Ridge regression, LoRA, or Skip-LoRA onto the corresponding 2D flat map of the NSD dataset. The results are presented in Figure 6. For linear heads, the importance of visual cortical voxels is similar for both ridge and LoRA. However, for Skip-LoRA, a small portion of the early visual cortex and more advanced visual cortex are valued more. This indicates that at the fMRI level, a greater concentration of non-linear relationships is found within the higher visual cortex. It also demonstrates that the design of Skip-LoRAs is capable of capturing non-linear relationships in fMRI data that are not discernible with standard LoRAs. More visualization results, as well as brain region ROI templates, can be found in Appendix E.
Figure 6: The importance of different ROIs of subject 1 with fitted weights of the first layer. The weights of each module are averaged and normalized between 0 and 1.
Conclusion
In this paper, we propose MindTuner, a new cross-subject decoding method. We introduced the phenomenon of visual fingerprint in the human visual system and utilized the combination of Skip-LoRAs and LoRAs to learn each subject’s visual fingerprint. Meanwhile, we innovatively propose a method for enhancing reconstruction by indirectly connecting fMRI with text in visual decoding tasks. Experimental results have shown that we have achieved better performance on multiple evaluation metrics at a relatively small parameter cost, especially when the fMRI data is insufficient. Our work has relaxed the conditions for fMRI acquisition, helping to achieve a universal brain decoding model in the future.
Acknowledgements
We have included more experimental results and analyses in the appendix, which you can access through this arXiv link [^2] https://arxiv.org/abs/2404.12630.
This research is supported by the National Key Research and Development Program of China (No. 2022YFB3104700), the National Natural Science Foundation of China (No. 61976158, No. 62376198), Shanghai Baiyulan Pujiang Project (No. 08002360429).